OpenStudy (anonymous):

Explain, in complete sentences, how you would divide the following expression: ^10√x^7 divided by ^7√x^4 Write the simplified answer in radical form

OpenStudy (anonymous):

\[\frac{\sqrt[10]{x^7}}{\sqrt[7]{x^4}}\]?

OpenStudy (anonymous):

YES !

OpenStudy (anonymous):

i would rewrite in exponential form, then subtract the exponent of the denominator from the exponent of the numerator, then rewrite in radical form

OpenStudy (anonymous):

that is the english, here is the math \[\large \frac{\sqrt[10]{x^7}}{\sqrt[7]{x^4}}=x^{\frac{7}{10}-\frac{4}{7}}\] \[\large x^{\frac{9}{70}}=\sqrt[70]{x^9}\]

OpenStudy (anonymous):

and that's the answer ?

OpenStudy (anonymous):

that is my answer, yes

OpenStudy (anonymous):

thank you so much !